# Nonparametric Estimation of Joint Entropy through Partitioned Sample-Spacing Method

### v1

@@ Abstract
We propose a nonparametric estimator of multivariate joint entropy based on partitioned sample spacings (PSS). The method extends univariate spacing ideas to $\mathbb{R}^{d}$ by partitioning into localized cells and aggregating within-cell statistics, with strong consistency guarantees under mild conditions. In benchmarks across diverse distributions, PSS consistently outperforms $k$-nearest neighbor estimators and achieves accuracy competitive with recent normalizing flow–based methods, while requiring no training or auxiliary density modeling. The estimator scales favorably in moderately high dimensions (d=10d=104040) and shows particular robustness to correlated or skewed distributions. These properties positions PSS as a practical alternative to NF-based approaches, with broad potential in information-theoretic machine learning applications.
00footnotetext: Corresponding author: hoo0321@yonsei.ac.kr 1Department of Statistics and Data Science, Yonsei University 2Department of Statistics and Probability, Michigan State University
@@ 1 Introduction
Entropy and mutual information are fundamental quantities in data analysis and machine learning, as they quantify uncertainty and dependence between variables Shannon1948. Unlike correlation, which captures only linear relationships, they can detect general nonlinear dependencies and remain invariant under monotone transformations. These properties have made entropy and mutual information central tools for feature selection, representation learning, and causal discovery.
@@
Estimating entropy and mutual information from samples is challenging. Among nonparametric approaches, the $k$-nearest neighbor (kNN) estimator is widely used PhysRevE.69.066138; Paninski20031191. It requires no distributional assumptions and adapts naturally to new data, but its accuracy degrades in multivariate settings, particularly under strong correlation or high dimension. More broadly, there is a long line of work on nonparametric estimation of divergences and conditional information using kkNN statistics poczos, which provides consistency guarantees but suffers from the same scalability issues.Recent refinements such as bias-corrected weighting BerrettSamworthYuan2019 improve asymptotic properties, yet all kNN variants rely on costly neighbor searches and remain sensitive to dependence.
@@
Another line of work leverages copula decompositions. Ariel and Louzoun copula proposed the Copula Decomposition Entropy Estimator (CADEE), which separates marginal and copula contributions using Sklar’s theorem. By recursively estimating copula entropy on the compact $0,1]^{d}$ domain, CADEE achieves strong robustness and scales better than kNN, particularly when supports are mixed or unbounded.
@@
In parallel, trainable normalizing flow methods have emerged. Ao and Li AoLi2022 introduced an estimator that learns a bijection to a tractable base distribution, while Butakov and Tolmachev ButakovTolmachev2024 proposed MIENF, which fits a pair of flows with finite-sample guarantees. Saad et al. saad developed inference-based entropy estimators (EEVI) that compute bounds within probabilistic generative models. While such estimators can reduce bias, they require GPU-scale training or access to tractable model densities, and may fail when the model family is misspecified.
@@
In this work, we revisit Vasicek’s classical sample-spacing idea Dudewicz1986161 and extend it to the multivariate setting by applying it locally within partitions of the sample space. The resulting estimator is fully nonparametric, avoids both neighbor search and neural training, and remains robust under heavy tails and nonlinear dependence. Under mild assumptions—boundedness and continuity of the density within each partition—we prove strong consistency, with an additional moment condition for entropy. Empirically, our method consistently outperforms kNN in mean-squared error, improves upon CADEE in correlated and moderately high-dimensional settings, and achieves accuracy competitive with normalizing-flow approaches while being substantially more efficient.
@@ 2.1 Nearest Neighbor–Based Estimators
where $\psi$ is the digamma function. Despite its simplicity, the KL estimator can suffer from large bias in high dimensions gao2015efficientestimationmutualinformation.
@@
Boundary bias can be severe when the support is compact. To address this, Ao and Li AoLi2022 introduced truncated versions (tKL, tKSG), where neighbor cells are clipped to the support $0,1]^{d}$. They proved that tKL is unbiased for uniform distributions. Building on this property, they proposed a normalizing flow transformation that maps samples approximately to uniform, and then applies the truncated estimator with a Jacobian correction. This yields the so-called uniformizing mapping (UM) estimators (UM-tKL, UM-tKSG), which achieve substantially lower bias in high dimensions. Normalizing flows themselves are neural network models of invertible transformations RezendeMohamed2015; Papamakarios2021. They require GPU-based training with many gradient updates, making them far more computationally demanding than nonparametric methods. Thus NF-based entropy estimation can reach high accuracy but at the cost of significant training overhead.
@@ 2.2 Univariate Case
Let $x_{(1)}<x_{(2)}<\cdots<x_{(n)}$ be order statistics of an i.i.d. sample $x_{1},x_{2},\dots,x_{n}$ drawn from a continuous random variable $X$ with probability density function $f(x)$. The difference $x_{(i+m)}-x_{(i)}$ is called m-spacing Learned-Miller20041271. Using this idea, park_and_park proposed the following density estimator based on sample spacings:
@@
This entropy estimator converges to $H(f)$ as $n,m\to\infty$ and $\frac{m}{n}\to 0$ Dudewicz1986161.
@@
The choice of the spacing parameter $m$ is critical for balancing bias and variance. To ensure consistency, $m$ must grow with the sample size nnBert-vanEs. For our theoretical results and empirical studies, we adopt the widely used rate $m=\lfloor\sqrt{n}+1/2\rfloor$ Crzc, which is known to be effective for smooth distributions.
Theorem 1 (Consistency of $\hat{f}_{n}(x)$).
If the spacing parameter is set as $m=\lfloor\sqrt{n}+1/2\rfloor$, then the sample-spacing density estimator $\hat{f}_{n}(x)$ converges almost surely to $f(x)$.
$\hat{f}_{n}(x)\xrightarrow{a.s.}f(x)$
@@
This result formally establishes the almost sure consistency of the univariate sample-spacing density estimator, providing a theoretical foundation for its extension to the multivariate and partitioned settings developed in the next sections.
@@ 3.1 Bivariate Case
Let $x_{(1)}\leq\dots\leq x_{(n)}$ and $y_{(1)}\leq\dots\leq y_{(n)}$ be the order statistics of $\{X_{i}\}$ and $\{Y_{i}\}$, respectively. Define the averaged grid points $\xi_{i}=\frac{x_{(i-m_{x})}+\cdots+x_{(i+m_{x}-1)}}{2m_{x}}$, $\eta_{j}=\frac{y_{(j-m_{y})}+\cdots+y_{(j+m_{y}-1)}}{2m_{y}}$.
with $x_{(i)}=x_{(1)}$ for $i<1$, $x_{(i)}=x_{(n)}$ for $i>n$, and analogously for $y_{(j)}$.
@@
Unlike adaptive partitioning approaches based on independence tests Darbellay19991315, or sample-spacing methods assuming radial symmetry Lee20102208, our method uses fixed, assumption-free partitions, making it both simple and broadly applicable.
@@ 3.2 The Partitioned Density Estimator
$\frac{n_{k}}{n}\frac{4m_{k}^{2}}{n_{k}^{2}(x_{(i+m_{k})}^{k}-x_{(i-m_{k})}^{k})(y_{(j+m_{k})}^{k}-y_{(j-m_{k})}^{k})}$
where $(x,y)\in(\xi_{i}^{k},\xi_{i+1}^{k}]\times(\eta_{j}^{k},\eta_{j+1}^{k}]$. The grid points are defined as $\xi_{i}^{k}=\frac{x_{(i-m_{k})}^{k}+\dots+x_{(i+m_{k}-1)}^{k}}{2m_{k}}$ and $\eta_{j}^{k}=\frac{y_{(j-m_{k})}^{k}+\dots+y_{(j+m_{k}-1)}^{k}}{2m_{k}}$, with boundary points $\xi_{0}^{k}=x_{(1)}^{k}$ and $\xi_{n_{k}+1}^{k}=x_{(n_{k})}^{k}$. Analogously, $\eta_{0}^{k}=y_{(1)}^{k}$ and $\eta_{n_{k}+1}^{k}=y_{(n_{k})}^{k}$. The partitioned density estimator is defined locally within each partition. To make the estimator more adaptive and reduce estimation bias, the sub-grid is constructed on the empirical support of the data within each partition, i.e., on the partition $\xi_{0}^{k},\xi_{n_{k}+1}^{k}]\times[\eta_{0}^{k},\eta_{n_{k}+1}^{k}]$. While this means the density is implicitly zero in the regions between the data’s range and the partition boundaries for finite samples, this choice is empirically superior, yielding a significantly lower MSE. As a result of this data-driven construction, the estimator is not guaranteed to integrate to exactly 1 for finite samples; however, it is asymptotically integrates to 1, as the empty regions between the data’s range and the partition boundaries vanish as $n_{k}\rightarrow\infty$.
Proposition 1.
@@
$\lim_{n\to\infty}\int_{\mathbb{R}^{2}}\hat{f}_{n,\ell}(x,y)\,dx\,dy=1$
This estimator is a valid probability density function and is strongly consistent. We formalize this in the following lemma and theorem. The proof strategy is to first establish, in lemma 2, the almost sure convergence of the univariate $m$-spacing estimators for the conditional distributions within each partition. This result serves as the core ingredient for Theorem 2, which combines these components to prove the consistency of the overall joint density estimator.
The assumptions in the lemma are critical for this two-stage argument. The conditions on $\ell(n)$ and $n_{k}$ ensure that as the total sample size $n$ grows, the partitions shrink appropriately while still accumulating enough data for asymptotic analysis. Similarly, the condition on the spacing parameter $m_{k}$ is crucial. As established in the preceding section, we adopt the specific rate $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$, which is required to prove convergence. Hence, we assume the conditions on $\ell(n)$, $n_{k}$, and the specific rate for $m_{k}$ holds for all subsequent asymptotic results.
Lemma 2.
Let the event $C_{k}\;=\;\bigl\{(X,Y)\in P_{k}\bigr\},$ $n_{k};=\sum_{v=1}^{n}\mathbf{1}\{(X_{v},Y_{v})\in C_{k}\}.$ Assume that as $n\to\infty$, $\ell(n)\to\infty,\ \ell(n)^{2}=o(n),\ n_{k}\rightarrow\infty.$ Further, assume the spacing parameter $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$. For $x\in(\xi_{j}^{k},\xi_{j+1}^{k}],\ y\in(\eta_{i}^{k},\eta_{i+1}^{k}]$ define the univariate mm‐spacing estimators
$\widehat{f}_{X,k}(x)=\frac{2\,m_{k}}{\,n_{k}\bigl(x_{k,(i+m_{k})}-x_{k,(i-m_{k})}\bigr)}$
@@
$\widehat{f}_{X,k}(x)\;\xrightarrow{a.s.}\;f_{X\mid C_{k}}(x),\ \widehat{f}_{Y,k}(y)\;\xrightarrow{a.s.}\;f_{Y\mid C_{k}}(y).$
Theorem 2 (Consistency of $\hat{f}_{n,\ell}(x,y)$).
@@
$\hat{f}_{n,\ell}(x,y)=\frac{n_{k}}{n}\cdot\hat{f}_{X;k}(x)\cdot\hat{f}_{Y;k}(y),$
@@
The almost sure convergence in Theorem 2 provides the foundation for $L^{1}$ convergence via Scheffé’s theorem Williams: if a sequence of nonnegative integrable functions $g_{n}$ satisfies (i) $g_{n}\to g$ a.e. and (ii) $\int g_{n}\to\int g$, then $\int|g_{n}-g|\to 0$. In our setting, Theorem 2 yields (i) for $\hat{f}_{n,\ell}$, and Proposition 1 establishes (ii) with $\int f=1$.
Corollary 1 ($L^{1}$ convergence of $\hat{f}_{n,\ell}$).
By Scheffé’s theorem, the L1 distance between the estimator and the true density converges to zero:
@@ 3.3 The Joint Entropy Estimator
Given the consistent density estimator $\hat{f}_{n,\ell}(x,y)$, a natural and powerful way to estimate the joint entropy is via a ”plug-in” approach. We replace the true density $f$ with our estimator $\hat{f}_{n,\ell}$ in the definition of entropy and approximate the expectation using the empirical distribution of the $n$ observed samples. This leads to the final entropy estimator:
@@
By applying the partition parameter $\ell$, several advantages arise. Since this estimator is computed locally, it is expected to perform well on highly correlated data and to effectively capture details of the distribution, such as skewness and kurtosis.
Additionally, the number of grids formed by the partitioned model is considerably smaller than in the previous model, making it more computationally efficient. For example, when $n$ samples are given, the previous model forms $n^{2}=\left(\sum_{k=1}^{\ell}n_{k}\right)^{2}$ grids, whereas the partitioned model forms $\sum_{k=1}^{\ell}n_{k}^{2}$ grids. As a result, the model becomes more computationally efficient as $\ell$ increases. However, factors such as variance, correlation, and sample size should also be considered when choosing the optimal value of $\ell$.
The optimal $\ell$ is expected to be invariant to variance, since the length of each partition is set as $\frac{\max(x)-\min(x)}{\ell}$ and $\frac{\max(y)-\min(y)}{\ell}$. Since these lengths are proportional to the variances of $X$ and $Y$, the number of samples $n_{k}$ in each partition is not influenced by variance for a fixed $\ell$. We can formally demonstrate this.
Proposition 2.
@@
Building upon the consistency of the density estimator, we now establish the $L^{1}$ convergence of the plug-in entropy estimator. The crucial step is to prove the uniform integrability of the log-density estimator sequence. This requires the globally adopted rate for the spacing parameter, $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$, as well as a mild integrability condition on the true density $f$ itself. The following theorem formalizes this main result.
Theorem 3.
Assume that the true density $f(x,y)$ satisfies $\log f\in L^{1+\delta}$ for some $\delta\in(0,1)$. Let the spacing parameter be $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$ with $n_{k}\geq 1$, and let the number of partitions satisfy $\ell(n)^{2}=o(n)$. Then, as $n\to\infty$
$\hat{H}_{n,\ell}(X,Y)\xrightarrow{L^{1}}H(f):=\mathbb{E}[-\log f(X,Y)]$

### v2

# Nonparametric Estimation of Joint Entropy via Partitioned Sample-Spacing

@@ Abstract
We propose a nonparametric estimator of multivariate joint entropy based on partitioned sample spacing (PSS). The method extends univariate spacing ideas to $\mathbb{R}^{d}$ by partitioning into localized cells and aggregating within-cell statistics, with strong consistency guarantees under mild conditions. In benchmarks across diverse distributions, PSS consistently outperforms $k$-nearest neighbor estimators and achieves accuracy competitive with recent normalizing flow–based methods, while requiring no training or auxiliary density modeling. The estimator scales favorably in moderately high dimensions (d=10d=104040) and shows particular robustness to correlated or skewed distributions. These properties position PSS as a practical and reliable alternative to both kkNN and NF-based entropy estimators, with broad utility in information-theoretic machine learning tasks such as total-correlation estimation, representation learning, and feature selection.
Keywords: mutual information, high-dimensional statistics, multivariate density estimation, information-theoretic machine learning
@@ 1 Introduction
Entropy and mutual information are fundamental quantities in data analysis and machine learning, as they quantify uncertainty and dependence between variables (Shannon1948). Unlike correlation, which captures only linear relationships, they can detect general nonlinear dependencies and remain invariant under monotone transformations. These properties have made entropy and mutual information central tools for feature selection, representation learning, independent component analysis and causal discovery.
@@
Estimating entropy and mutual information from samples is challenging. Among nonparametric approaches, the $k$-nearest neighbor (kNN) estimator is widely used (kraskov2004). It requires no distributional assumptions and adapts naturally to new data, but its accuracy degrades in multivariate settings, particularly under strong correlation or high dimension. More broadly, there is a long line of work on nonparametric estimation of divergences and conditional information using kkNN statistics (poczos), which provides consistency guarantees but suffers from the same scalability issues. Recent refinements such as bias-corrected weighting (BerrettSamworthYuan2019) improve asymptotic properties, yet all kNN variants rely on costly neighbor searches and remain sensitive to dependence.
@@
Another line of work leverages copula decompositions. Ariel and Louzoun (copula) proposed the Copula Decomposition Entropy Estimator (CADEE), which separates marginal and copula contributions using Sklar’s theorem. By recursively estimating copula entropy on the compact $0,1]^{d}$ domain, CADEE achieves strong robustness and scales better than kNN, particularly when supports are mixed or unbounded.
@@
In parallel, trainable normalizing flow methods have emerged. AoLi2022 introduced an estimator that learns a bijection to a tractable base distribution, while ButakovTolmachev2024 proposed MIENF, which fits a pair of flows with finite-sample guarantees. saad developed inference-based entropy estimators (EEVI) that compute bounds within probabilistic generative models. While such estimators can reduce bias, they require significant computational resources for training or access to tractable model densities, and fail when the model family is misspecified.
@@
In this work, we revisit Vasicek’s classical sample-spacing idea (Vasicek)(Dudewicz1986161) and extend it to the multivariate setting by applying it locally within partitions of the sample space. The resulting estimator is fully nonparametric, avoids both neighbor search and model training, and remains robust under heavy tails and nonlinear dependence. Under mild assumptions—boundedness and continuity of the density within each partition—we prove strong consistency, with an additional moment condition for entropy. Empirically, our method consistently outperforms kNN in mean-squared error, improves upon CADEE in correlated and moderately high-dimensional settings, and achieves accuracy competitive with normalizing-flow approaches while being substantially more efficient. These results position partitioned sample spacings as a practical and data-efficient tool for modern information-theoretic learning tasks.
@@ 2.1 Nearest Neighbor–Based Estimators
where $\psi$ is the digamma function. Despite its simplicity, the KL estimator can suffer from large bias in high dimensions (gao2015efficientestimationmutualinformation).
@@
Boundary bias can be severe when the support is compact. To address this, AoLi2022 introduced truncated versions (tKL, tKSG), where neighbor cells are clipped to the support $0,1]^{d}$. They proved that tKL is unbiased for uniform distributions. Building on this property, they proposed a normalizing flow transformation that maps samples approximately to uniform, and then applies the truncated estimator with a Jacobian correction. This yields the so-called uniformizing mapping (UM) estimators (UM-tKL, UM-tKSG), which achieve substantially lower bias in high dimensions. Normalizing flows themselves are neural network models of invertible transformations (RezendeMohamed2015; Papamakarios2021). They require GPU-based training with many gradient updates, making them far more computationally demanding than nonparametric methods. Thus NF-based entropy estimation can reach high accuracy but at the cost of significant training overhead.
@@ 2.2 Univariate Case
Let $x_{(1)}<x_{(2)}<\cdots<x_{(n)}$ be order statistics of an i.i.d. sample $x_{1},x_{2},\dots,x_{n}$ drawn from a continuous random variable $X$ with probability density function $f(x)$. The difference $x_{(i+m)}-x_{(i)}$ is called m-spacing (Learned-Miller20041271). Using this idea, park_and_park proposed the following density estimator based on sample spacings:
@@
This entropy estimator converges to $H(f)$ as $n,m\to\infty$ and $\frac{m}{n}\to 0$ (Dudewicz1986161).
@@
See Appendix A for the proof. The choice of the spacing parameter $m$ is critical for balancing bias and variance. To ensure consistency, $m$ must grow with the sample size $n$ (Bert-vanEs). For our theoretical results and empirical studies, we adopt the widely used rate $m=\lfloor\sqrt{n}+1/2\rfloor$ (Crzc), which is known to be effective for smooth distributions.
Theorem 2 (Consistency of $\hat{f}_{n}(x)$)
If the spacing parameter is set as $m=\lfloor\sqrt{n}+1/2\rfloor$, then the sample-spacing density estimator $\hat{f}_{n}(x)$ converges almost surely to $f(x)$:
$\hat{f}_{n}(x)\xrightarrow{a.s.}f(x),$
@@
See Appendix B for the proof. This result formally establishes the almost sure consistency of the univariate sample-spacing density estimator, providing a theoretical foundation for its extension to the multivariate and partitioned settings developed in the next sections.
@@ 3.1 Bivariate Case
Let $x_{(1)}\leq\dots\leq x_{(n)}$ and $y_{(1)}\leq\dots\leq y_{(n)}$ be the order statistics of $\{X_{i}\}$ and $\{Y_{i}\}$, respectively. Define the averaged grid points $\xi_{i}=\frac{x_{(i-m_{x})}+\cdots+x_{(i+m_{x}-1)}}{2m_{x}}$, $\eta_{j}=\frac{y_{(j-m_{y})}+\cdots+y_{(j+m_{y}-1)}}{2m_{y}}$, with $x_{(i)}=x_{(1)}$ for $i<1$, $x_{(i)}=x_{(n)}$ for $i>n$, and analogously for $y_{(j)}$.
@@
Unlike adaptive partitioning approaches based on independence tests (Darbellay19991315), or sample-spacing methods assuming radial symmetry (Lee20102208), our method uses fixed, assumption-free partitions, making it both simple and broadly applicable.
@@ 3.2 The Partitioned Density Estimator
$\hat{f}_{n,\ell}(x,y)=\frac{n_{k}}{n}\frac{4m_{k}^{2}}{n_{k}^{2}(x_{(i+m_{k})}^{k}-x_{(i-m_{k})}^{k})(y_{(j+m_{k})}^{k}-y_{(j-m_{k})}^{k})}$
where $(x,y)\in(\xi_{i}^{k},\xi_{i+1}^{k}]\times(\eta_{j}^{k},\eta_{j+1}^{k}]$. The grid points are defined as $\xi_{i}^{k}=\frac{x_{(i-m_{k})}^{k}+\dots+x_{(i+m_{k}-1)}^{k}}{2m_{k}}$ and $\eta_{j}^{k}=\frac{y_{(j-m_{k})}^{k}+\dots+y_{(j+m_{k}-1)}^{k}}{2m_{k}}$, with boundary points $\xi_{0}^{k}=x_{(1)}^{k}$ and $\xi_{n_{k}+1}^{k}=x_{(n_{k})}^{k}$. Analogously, $\eta_{0}^{k}=y_{(1)}^{k}$ and $\eta_{n_{k}+1}^{k}=y_{(n_{k})}^{k}$. The partitioned density estimator is defined locally within each partition.
To make the estimator more adaptive and reduce estimation bias, the sub-grid is constructed on the empirical support of the data within each partition, i.e., on the partition $\xi_{0}^{k},\xi_{n_{k}+1}^{k}]\times[\eta_{0}^{k},\eta_{n_{k}+1}^{k}]$. While this means the density is implicitly zero in the regions between the data’s range and the partition boundaries for finite samples, this choice is empirically superior, yielding a significantly lower MSE. As a result of this data-driven construction, the estimator is not guaranteed to integrate to exactly 1 for finite samples; however, it asymptotically integrates to 1, as the empty regions between the data’s range and the partition boundaries vanish as $n_{k}\rightarrow\infty$.
Proposition 3
@@
$\lim_{n\to\infty}\int_{\mathbb{R}^{2}}\hat{f}_{n,\ell}(x,y)\,dx\,dy=1.$
See Appendix C for the proof. This estimator is a valid probability density function and is strongly consistent. We formalize this in the following lemma and theorem. The proof strategy is to first establish, in Lemma 4, the almost sure convergence of the univariate $m$-spacing estimators for the conditional distributions within each partition. This result serves as the core ingredient for Theorem 5, which combines these components to prove the consistency of the overall joint density estimator.
The assumptions in the lemma are critical for this two-stage argument. The conditions on $\ell(n)$ and $n_{k}$ ensure that as the total sample size $n$ grows, the partitions shrink appropriately while still accumulating enough data for asymptotic analysis. Similarly, the condition on the spacing parameter $m_{k}$ is crucial. As established in the preceding section, we adopt the specific rate $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$, which is required to prove convergence. Hence, we assume the conditions on $\ell(n)$, $n_{k}$, and the specific rate for $m_{k}$ hold for all subsequent asymptotic results.
Lemma 4
Let the event $C_{k}\;=\;\bigl\{(X,Y)\in P_{k}\bigr\}$ and $n_{k}:=\sum_{v=1}^{n}\mathbf{1}\{(X_{v},Y_{v})\in C_{k}\}.$ Assume that as $n\to\infty$, $\ell(n)\to\infty,\ \ell(n)^{2}=o(n),\ n_{k}\rightarrow\infty.$ Further, assume the spacing parameter $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$. For $x\in(\xi_{j}^{k},\xi_{j+1}^{k}],\ y\in(\eta_{i}^{k},\eta_{i+1}^{k}]$ define the univariate mm‐spacing estimators
$\widehat{f}_{X,k}(x)$ $=\frac{2\,m_{k}}{\,n_{k}\bigl(x_{k,(i+m_{k})}-x_{k,(i-m_{k})}\bigr)},$
@@
$\widehat{f}_{X,k}(x)\;\xrightarrow{a.s.}\;f_{X\mid C_{k}}(x),\quad\widehat{f}_{Y,k}(y)\;\xrightarrow{a.s.}\;f_{Y\mid C_{k}}(y).$
See Appendix D for the proof.
Theorem 5 (Consistency of $\hat{f}_{n,\ell}(x,y)$)
@@
$\hat{f}_{n,\ell}(x,y)=\frac{n_{k}}{n}\cdot\widehat{f}_{X,k}(x)\cdot\widehat{f}_{Y,k}(y),$
@@
See Appendix E for the proof. The almost sure convergence in Theorem 5 provides the foundation for $L^{1}$ convergence via Scheffé’s theorem (Williams): if a sequence of nonnegative integrable functions $g_{n}$ satisfies (i) $g_{n}\to g$ a.e. and (ii) $\int g_{n}\to\int g$, then $\int|g_{n}-g|\to 0$. In our setting, Theorem 5 yields (i) for $\hat{f}_{n,\ell}$, and Proposition 1 establishes (ii) with $\int f=1$.
Corollary 6 ($L^{1}$ convergence of $\hat{f}_{n,\ell}$)
By Scheffé’s theorem, the $L^{1}$ distance between the estimator and the true density converges to zero:
@@ 3.3 The Joint Entropy Estimator
Given the consistent density estimator $\hat{f}_{n,\ell}(x,y)$, a natural and powerful way to estimate the joint entropy is via a “plug-in” approach. We replace the true density $f$ with our estimator $\hat{f}_{n,\ell}$ in the definition of entropy and approximate the expectation using the empirical distribution of the $n$ observed samples. This leads to the final entropy estimator:
@@
Applying the partition parameter $\ell$ offers several advantages. Since this estimator is computed locally, it adapts well to highly correlated data and effectively captures distributional details such as skewness and kurtosis.
Additionally, the partitioned model significantly improves computational efficiency by reducing the number of effective grid cells. For example, with $n$ samples, the global (unpartitioned) method implicitly forms a grid of size $n^{2}$, whereas the partitioned model constructs independent grids summing to $\sum_{k=1}^{\ell}n_{k}^{2}$. Since $\sum n_{k}^{2}\ll(\sum n_{k})^{2}$, the computational burden decreases as $\ell$ increases. However, the choice of $\ell$ must also balance the trade-off involving variance, correlation, and sample size.
Notably, the optimal $\ell$ is expected to be invariant to the scale of the data. Because the length of each partition is set as proportional to the range (e.g., $\frac{\max(x)-\min(x)}{\ell}$), the number of samples $n_{k}$ falling into each partition remains constant under affine transformations. We formally demonstrate this property below.
Proposition 7
@@
See Appendix F for the proof. Building upon the consistency of the density estimator, we now both establish the almost sure and the $L^{1}$ convergence of the plug-in entropy estimator. The crucial step is to prove the uniform integrability of the log-density estimator sequence. This requires the globally adopted rate for the spacing parameter, $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$, as well as a mild integrability condition on the true density $f$ itself. The following theorem formalizes this main result.
Theorem 8
Assume that the true density $f(x,y)$ satisfies $\log f\in L^{1+\delta}$ for some $\delta\in(0,1)$. Let the spacing parameter be $m_{k}=\lfloor\sqrt{n_{k}}+1/2\rfloor$ with $n_{k}>1$, and let the number of partitions satisfy $\ell(n)^{2}=o(n)$. Then, as $n\to\infty$, the entropy estimator $\hat{H}_{n,\ell}(X,Y)$ converges to the true entropy $H(f):=\mathbb{E}[-\log f(X,Y)]$ both almost surely and in $L^{1}$:
$\hat{H}_{n,\ell}(X,Y)\xrightarrow{a.s.}H(f)\quad\text{and}\quad\hat{H}_{n,\ell}(X,Y)\xrightarrow{L^{1}}H(f).$
See Appendix G for the proof.